Solve by using the quadractic formula.
x^2=7x+6

Answer by stanbon(75887)

Solve by using the quadractic formula.
x^2=7x+6
x^2-7x-6=0
x = [7 +- sqrt(49-4*-6)]/2
x = [7 +- sqrt(73)]/2
x = (7+sqrt73)/2 or x = (7-sqrt73)/2
=============
Cheers,
Stan H.